# Multi-Type Point Cloud Autoencoder: A Complete Equivariant Embedding for Molecule Conformation and Pose

### v1

@@ 3.4 Distance Calculation
Beyond the reconstruction loss, which is somewhat abstract in chemical structure terms, we calculate the root mean squared distance (RMSD) between input and output point clouds. To accomplish this, the output points are first clustered into predicted particle types and positions, to match against the inputs. When given a molecular scaffold, this can be straightforwardly done by assigning any output cloud points $j$ within less than half a bond length of inputs $i$ to cluster $i$, and aggregating them by their respective weights. In the case of no molecular scaffold, such as during de novo generation or conformational transformation, agglomerative clustering is performed on the output swarm in the Cartesian dimensions, followed by a merge of low-weight clusters into their nearest neighbors.
@@
Both of these approaches qualitatively agree with visual inspection and the reconstruction loss in the range of small deviations, where the output swarm particles from a trained autoencoder model are generally tightly clustered around the predicted particle positions. When the output swarm is diffuse, meaning the reconstruction is under-converged, the reconstruction loss is large, and both clustering approaches fail. In other words, these clustering approaches only return sensible results when applied to an already reasonably good reconstruction.
3.5 Training protocol
During training, when the loss reaches a certain lower threshold, we gradually decrease $\sigma$ such that eq. (4) converges to eq. (3). This gradually increases the difficulty of the problem, as the target distribution narrows, until the mean self-overlap of the input samples is below 0.0005. All model hyperparameters and further training details are given in the supplemental material.
Some molecular datasets are published with absent, fuzzy, or incomplete hydrogen information. It is therefore useful to have embedding models appropriate for either the presence or absence of hydrogens. Thanks to the flexibility of our approach, we can easily test the model’s ability to accurately place hydrogen atoms given an input where hydrogen atoms have been removed.
@@ 4.1 Autoencoder performance
We begin by assessing the reconstruction performance of the Mo3ENet autoencoder on the QM9 dataset under three conditions: (1) no hydrogens in input or output, and RMSD calculated on heavy atoms only, (2) hydrogens in input and output, with RMSD calculated on all atoms, and (3), no hydrogens in input, but hydrogen positions inferred in output, with RMSD calculated on all atoms.
Figure 4: RMSD distribution for three autoencoder models on the full QM9 dataset. The models omit, reconstruct, and infer the locations of hydrogen atoms, respectively. Fraction of molecules which where all atoms could not be uniquely matched are noted for each model.
@@
Training results are summarized in Figure 4, where we show the distribution of RMSDs for the three cases, for the samples where input and input atoms could be uniquely matched by our scaffolding algorithm. All three models converge well with mean RMSDs $\leq 0.1$Å, and with a reasonably tight distribution about the mean. We note for each model the mean deviation and fraction of the dataset for which the scaffolding failed to uniquely match inputs to outputs. These generally correspond to cases where either the output swarm was too diffuse to be neatly clustered, or where one or more atoms were simply not properly captured, even if the rest of the molecule is correct.
A significant positive correlate of error is the fraction of fluorine in each molecule, which should, perhaps, not be surprising since fluorine is by far the rarest atom type in the dataset. Possibly, even relatively coarse sample augmentation such as substituting hydrogens for fluorines in some samples could improve generalization in this dimension, though we avoided such tinkering in this study. All models overfit the reconstruction loss to some degree, and so it seems likely that improved performance could be achieved on larger datasets, or with additional hyperparameter tuning.
@@
Figure 5: The UMAP decomposition for the full QM9 dataset according to our autoencoder trained with hydrogens, with the x and y axes the UMAP reduced dimensions of the 512 dimensional encoder embedding. Points are color coded according to compositional (atom fraction) and geometric molecular (principal inertial ratios) factors, with legends for each on the right hand side.
@@
In Figure 5, we map the latent space of the autoencoder embedding in a reduced dimensional space. There is an open choice of which form of the embedding to visualize here, the full vector embedding, the scalarized embedding which includes a learned angular projection, or simply the Cartesian norms of the vector embedding. We chose the scalarized output (details of ‘scalarizer’ in SM), since it should contain richer information than the raw vector norms, including primarily the relative directions of vector components, and is invariant to molecule pose, which could over-complicate the mapping.
@@
We employ UMAP [57], a dimensionality reduction algorithm, similar in spirit to principal component analysis [58], or t-distributed stochastic neighborhood embedding [59], which attempts to preserve relative distances between samples, such that samples which are near or far in the high dimensional latent space are similarly placed in the reduced dimension. For visualization purposes, we generated 2-component projections from the scalarized embedding vectors of the full QM9 dataset, shown in Figure 5(a)-(b) together with various molecules noted in the margins, and colors assigned according to (a) relative atomic compositions and (b) shapes.
@@ 4.2 Property prediction
Figure 6: Predicted vs. true values on the test dataset (20% of sample) for the 15 QM9 quantum mechanical scalar properties, with associated correlation coefficients and mean absolute errors.
With a trained encoder in hand, we proceed to scalar molecular property prediction. We freeze the encoder weights and train separate EMLP models for regression of the 15 QM9 properties, directly on the molecule embeddings, omitting the decoder. Regression hyperparameters and training details are given in the SM. We found that the overall performance was relatively insensitive to the details of the regression head, and use the same architecture for all properties.
@@
In Figure 6, the regression results on the 15 molecular properties of the QM9 database are presented. Although our results do not surpass prior works with exhaustively pretrained encodings [18], errors are small across the board, indicating that the learned embedding is both complete in the sense of containing all the relevant molecular information, and well enough regularized that models trained on top of it are able to generalize well to the test data.
Figure 7: Training curve for regression of the HOMO-LUMO gap, showing the test loss minimum for each training run, as a function of total dataset size.
@@
A caveat to this is, as we see in Figure 7, dataset size is clearly limiting generalization performance, and we observed significant overfitting in training. This is likely due to both the architecture of the regression head and complexity of the embedding.
@@ 5 Discussion
We introduce and evaluate a new autoencoder architecture and reconstruction loss for the unsupervised learning of multi-type point clouds. This tool ‘solves’ the molecule representation problem in the limit of large data, in the sense that it can encode type, conformation, and pose information of a molecule in a provably complete embedding, via a sufficiently large and well trained autoencoder model. The main practical limitations of such a model are governed by the ability to efficiently optimize its parameters as well as the size and composition of the training dataset (equilibrium vs. non-equilibrium structures, molecule composition and size, etc.). The embedding quality on any new samples can, however, be trivially evaluated by decoding the sample and computing the reconstruction loss.
Beyond high-fidelity reconstructions, we further proved the quality of the molecule embedding as an input for regressing the fifteen QM9 molecular properties. The good performance on test data indicates that the embedding is to a certain extent physically meaningful, in contrast to a model that simply memorizes the dataset in a very large latent dimension. This point is buttressed by our embedding analyses, where we show visually that the model is clustering functional groups and structural motifs in its embedding space.
This work opens the door to several interesting applications. First, via training on larger and more diverse datasets, we can develop a universal/foundational representation model for high-fidelity encoding of molecule and pose information with this approach. Such a tool could be useful for molecule property prediction, similarity analysis/clustering, and conditioning for downstream learning tasks. The speed and generality of our approach also make it attractive as a generative model. The freedom of the decoder to generate structures with arbitrary numbers, types, and positions of atoms, and the latent space’s amenability to conditioning, is tantalizing to explore further. Finally, such a pretrained molecule embedding will be very useful in future exploration of molecule clusters and materials.

### v3

@@ 3.3 Distance Calculation
Beyond the reconstruction loss, which is somewhat abstract in chemical structure terms, we also calculate a mean distance between input and output point clouds, defined as distance between all input atoms and their partners in the decoded output (see the rigorous definition in the appendix).
@@
Since Mo3ENet is equivariant and molecules are always centered on the origin, there is no need for overall structure matching, only assignment of output GM components to input atoms. When given a molecular scaffold, this can be straightforwardly done by assigning any output cloud points $j$ within a short range (0.5Å) of inputs $i$ to cluster $i$, and aggregating them to synthetic particle predictions by averaging according to their respective weights. The atomwise deviations are then easily calculated and averaged molecule-wise. In the case of no molecular scaffold, such as during de novo generation or conformational transformation, agglomerative clustering can be performed on the output swarm in the Cartesian dimensions, followed by a merge of low-weight clusters into their nearest neighbors.
@@
Both of these approaches qualitatively agree with visual inspection and the reconstruction loss (probability mass overlap between input and output GMs) in the range of small deviations, where the output swarm particles from a trained autoencoder model are generally tightly clustered around the predicted particle positions. When the output swarm is diffuse, meaning the reconstruction is under-converged, the reconstruction loss is large, and both clustering approaches fail. In other words, these clustering approaches only return sensible results when applied to an already reasonably good reconstruction.
3.4 Molecular Data
The point clouds contained in molecular datasets such as QM9 are frequently optimized against some potential energy function, representing a ground state or zero temperature structure. While these conformers are ‘probable’ structures, molecules are generally at least somewhat flexible, and a given molecule may adopt a wide range of conformations. An autoencoder trained on only highly optimized structures may struggle to apprehend valid structures in non-ideal conformations, or conformations which are unlikely under a given potential. Therefore, to ensure generalization to a variety of molecular conformations, we developed a pipeline for synthesizing large numbers of random conformers.
Before training, we extracted all unique canonicalized SMILES codes from the QM9 and ZINC22 datasets containing H, C, N, O, and F, up to 9 and 14 heavy atoms, respectively. 10k optimized molecular structures from QM9 were set aside for evaluation, and the remaining 7.2M SMILES are used for training.
Point clouds are generated according to the following procedure: (1) select a random SMILES code from the training set, (2) generate a rough 3D conformation with RDKit [62], (3) if the molecule has more than 9 heavy atoms, iteratively truncate it at rotatable bonds, capping with hydrogens, until it is within the appropriate range, (4) randomly spin all rotatable bonds in the molecule, (5) run very short energy minimizations to avoid severe interatomic overlaps, (6) to further boost generalization to different molecule conditions, gently noise and globally rescale the completed structure.
This dataset synthesis approach provides a very wide range of molecular structures, practically limited only by the distribution of functional groups represented in the QM9+ZINC22 SMILES, which could be amended in the future with more datasets such as GEOM [63], or QMugs [64], or by adding random generation of the molecule backbone to the above procedure.
We also attempted training on point clouds sampled from a random uniform distribution in coordinates and atom types, but found negligible benefit to model performance on molecular data. This should, perhaps, not be surprising, since molecular information is structured by certain correlations, such as bond lengths and angles, which random point clouds lack, and therefore are far outside the relevant distribution.
@@ 4.1 Molecule reconstruction
We begin by assessing the reconstruction performance of the Mo3ENet autoencoder under two conditions: (1) no hydrogens in input or output, and calculated on heavy atoms only, and (2) hydrogens in input and output, with mean distances calculated on all atoms. Some molecular datasets are published with absent, fuzzy, or incomplete hydrogen information. It is therefore useful to have embedding models appropriate for either the presence or absence of hydrogens.
Figure 4: Mean distance distribution for Mo3ENet models on the test QM9 dataset (10k samples). The models omit or include hydrogen atoms, respectively. In panel (a) we show the mean deviations averaged over the whole molecule, for molecules where every atom was matched by our clustering procedure. We also give the overall average and matched molecule fractions. In panel (b) we show the same statistics on a per-atom basis.
@@
The reconstruction performance is summarized in Figure 4, where we show the distribution of atom-wise and molecule-averaged mean distances on the test dataset, for the samples where input and output atoms could be uniquely matched by our scaffolding algorithm. Both models converge well with molecular mean distances $<0.1$Å, and with a reasonably tight distribution about the mean. We note for each model the mean deviation and fraction of the dataset for which the scaffolding failed to uniquely match inputs to outputs. These generally correspond to cases where either the output swarm was too diffuse to be neatly clustered, or where one or more atoms were not properly captured, even if the rest of the molecule is correct. This interpretation is supported by the fact that the atom-by-atom matching fraction is significantly better than that for whole-graphs, indicating that molecules which fail to match are mostly complete, with a relatively small number of imperfectly captured atoms.
A significant positive correlate of error is the fraction of fluorine in each molecule, which should, perhaps, not be surprising since fluorine is by far the rarest atom type in the dataset. A wider range of molecular backbones would therefore likely boost overall generalization. Indeed, our experiments indicate longer training runs with more molecules and larger models should yield systematic improvements.
A second and interesting source of significant errors during development were highly symmetric molecules. We found that molecules with mirror planes or inversion centres frequently had extremely poor, sometimes practically random, reconstructions. This effect arises ironically from one of the design advantages of the model: equivariance. The final step in molecule encoding is aggregating atom-wise information to a molecule-scale descriptor, the molecule embedding. Whether the graph aggregation operation is based on a sum, mean, self-attention, or combination of the above, it requires summing of atom embedding vectors. In an equivariant model, if two atoms have identical local environments and are located opposite one another by a inversion centre, they will have identical node vectors, up to a global rotation; when summed, they exactly cancel, giving the decoder downstream no information at all to reconstruct the molecule.
Fortunately, we found empirically this could be ameliorated by injecting a tiny amount of noise into the atom positions (<0.01Å on average), lifting the offending symmetry and allowing the encoder to pass nonzero information to the decoder. Many of the lowest quality reconstructions, even after this patch, are still high-symmetry molecules, indicating it is still more difficult to learn their structures, but their reconstructions are still generally good and mean distances are within the normal range of the data.
4.2 Embedding analysis
@@
Figure 5: The UMAP decomposition for the full QM9 dataset according to our autoencoder trained with hydrogens, with the x and y axes the UMAP reduced dimensions of the 64 dimensional encoder embedding. Points are color coded according to, (a) compositional (atom fraction) and, (b) geometric molecular (principal inertial ratios) factors, with legends for each on the right hand side.
@@
In Figure 5, we map the encoder embedding in a reduced dimensional space. There is an open choice of which form of the embedding to visualize here, the full vector embedding, the scalarized embedding which includes a learned angular projection, or simply the Cartesian norms of the vector embedding. We chose the scalarized output (details of ‘scalarizer’ in the appendix), as it should contain richer information than the vector norms, including primarily the relative directions of vector components, and is invariant to molecule pose, which could over-complicate the mapping.
@@
We employ UMAP [65], a dimensionality reduction algorithm, similar in spirit to principal component analysis [66], or t-distributed stochastic neighborhood embedding [67], which attempts to preserve relative distances between samples, such that samples that are near or far in the high dimensional latent space are similarly placed in the reduced dimension. For visualization purposes, we generated 2-component projections from the scalarized embedding vectors of the full QM9 dataset, shown in Figure 5(a)-(b), with colors assigned according to (a) relative atomic compositions and (b) shapes. In the margins, we visualize molecules connected by arrows to their respective points in the embedding space, to give a qualitative sense of the types of molecules clustered in each region.
@@ 4.3 Property prediction
Figure 6: Predicted vs. true values on the test dataset (20% of sample) for the 15 QM9 scalar properties, and 5 QM9s vector and tensor properties, with associated correlation coefficients and mean absolute errors. Thermodynamic quantities are evaluated at 298K unless otherwise stated. Detailed plots for each property are provided in the appendix.
We have proven above that our molecule representation is complete and generalizes well to unseen QM9-like conformers. Given that the molecular embedding generated by the encoder is rich enough to reconstruct the molecule in full, it should serve as a good representation from which to predict molecule properties. Therefore, to demonstrate further its utility and generalizability, and begin to explore the Mo3ENet’s behavior on downstream modelling tasks, we undertake, below, scalar, vector, and tensor prediction tasks on 20 molecular properties.
For the prediction of the 15 scalar properties of the QM9 molecules, we trained separate multilayer perceptrons as regression heads on a frozen pretrained encoder. For the prediction of the 5 QM9s vector and tensor molecule properties [68], we trained equivariant multilayer perceptron regression models in the same way. QM9s re-optimizes the QM9 molecular geometries, adding dipole and multipole moments, and polaraizabilities. The dataset was split 80:20 into train:test to assess generalization to unseen molecules. Regression hyperparameters and training details are given in the appendix. We undertook a small-scale hyperparameter sweep and selected values which gave good test losses on the molecule gap energy, and reused them for all other properties.
We present the regression results in Figure 6, where we see generally very good fits across the board. More detailed visualizations are given in the appendix. HOMO, LUMO and gap energies are typically used for benchmarking on this dataset, and with an average MAE 0.00552, we are within a comparable range to the current benchmark value of 0.00467, set by the extensively pre-trained UniMol [13]. Vector/tensor property prediction is also qualitatively good, although we do only outperform state-of-the-art purpose-built models such as DetaNet [68] on one task, as shown in Table 1.
Table 1: MAE comparison on QM9s properties between Mo3ENet and DetaNet.
Model Dipole Polarizability Hyperpolarizability Quadrupole Octupole
Mo3ENet 0.173 2.32 1.25 14.9 11.1
DetaNet 0.0089 0.137 2.04 0.38 6.29
@@
For all 20 properties, the target:prediction correlations are high, indicating that the learned embedding is both complete in the sense of containing the relevant molecular information, and well enough regularized that models trained on top of it are able to generalize well to the test data.
@@
We see in our experiments that dataset size is clearly limiting generalization performance, and we observed significant overfitting in training on all properties. This is likely due to combination of the difficulty of the regression tasks, the architecture of the regression head, the complexity of the embedding, and imperfections in the Mo3ENet embeddings.
4.4 Crystal embedding model
Going one step further downstream, we introduce here a new learning task, taking advantage of the properties of the Mo3ENet encoding. Given the inherent relatively low-dimensionality of the parameter space of molecular crystals, which includes cell lengths, cell angles, molecular orientation, internal molecular conformation, molecular center-of-geometry location, and space group, we can combine a molecule embedding with only a few such crystal parameters to create an informationally complete representation, and train models in this space. We call this representation the crystal embedding,
$\mathcal{C}_{m}=m||\mathcal{C},$ (7)
for $m$ a molecule embedding or representation, and $\mathcal{C}$ the crystal parameters corresponding to the unit cell box vectors and and molecule placement in the asymmetric unit (see details in appendix).
As a crystal embedding with a sufficiently rich molecule embedding could be used in principle to model any crystal property. We selected the crystal intermolecular potential energy for this first task, as it is (1) relatively common and well-understood, and (2) quite physically rich. Molecular crystal energies depend in very complex ways on the details of intra- and inter-molecular structure, with, for example, a global rotation of a molecule of only a few degrees potentially resulting in severe interatomic clashes. Going further, the types of intermolecular interactions a molecule will participate in depend on its composition (presence of relevant functional groups) and conformation. Highly accurate prediction of crystal energies will therefore only be possible with a molecule embedding incorporating molecular conformation, composition, and pose, such as that from Mo3ENet.
We trained MLP regression heads to predict two types of qualitatively different intermolecular potentials, and compared the autoencoder embedding to a series of simpler molecule embeddings, to observe trends in model performance with increasing representational detail.
The first intermolecular energy function is a Buckingham potential,
$E_{BH}=\sum_{ij}A\cdot e^{-Bd_{ij}}-\frac{C}{d_{ij}^{6}},$ (8)
[…]
The second crystal energy for this benchmark is a MACE machine learned interatomic potential (MLIP) for atomistic systems [69], specifically the ‘medium-sized’ MACE-MPA-0 model, targeted at materials. Extensively pretrained, MACE-type models have shown reasonable performance on molecular and materials data. We use this model here not to compute highly accurate energies for molecular crystals but to capture far more nuanced and complex intermolecular interactions compared to a simple interatomic pair potential. To calculate the crystal lattice energy within a rigid molecule approximation, we compute separately the MACE-MPA-0 energy for a molecular crystal and gas phase molecule (in the crystal conformation without relaxation), and take the difference,
$E_{lattice}=\frac{E_{crystal}}{Z}-E_{gas},$ (9)
with $Z$ the number of molecules per unit cell.
We test four molecule embedding schemes, (1) an empty embedding, to challenge what the model can do with only the crystal lattice information (2) the molecule volume, to give the model some insight into crystal densities, (3) principal inertial vectors (eigenvectors and moments of the inertial tensor, equivalent to an ellipsoid description), which has been shown to be a good descriptor of crystal planes [70], and (4) Mo3ENet encoding. These embeddings are combined with the 12 crystal parameters for a rigid, $Z^{\prime}=1$ molecular crystal, $\mathcal{C}$ (details in the appendix), to form a representation for a crystal.
A dataset of $\sim$2.4M crystals in the $P$1 space group, comprising $\sim$220k unique molecules was synthesized and embedded according to these four schemes, and used to train an MLP (hyperparameters in the appendix) to predict the crystal energy. An additional 63k structures were synthesized for validation.
Crystal sample generation comprises the following steps: (1) generate a random conformer, (2) sample random crystal parameters and build a supercell of the relevant crystal (see SM for details), (3) optimize the crystal parameters against a Lennard-Jones energy function, holding the intramolecular degrees of freedom fixed, and add the final structure to the dataset. For additional diversity, we add 10 non-optimal samples per optimized crystal, each generated by adding a small amount of Gaussian noise to the optimized cell parameters. The resulting dataset covers a wide range of chemical structures, molecular conformations, and crystal topologies.
Figure 7: Parity plots for the crystal embedding energy regression task on the validation set, for several molecule embeddings and two energy functions. Subplots (a-d) show the regression results on a Buckingham potential, and (e-h) the same on the MACE MPA-0 potential.
We observe in Figure 7 sensible trends in the crystal embedding regression task. The relatively basic Buckingham energies are possible to estimate, at least somewhat accurately, given only knowledge of molecule size and unit cell volume, assuming minimal intermolecular atomic clashes, a reasonable assumption for this dataset, given its construction procedure. This is evidenced in subplot (b), where we already see that merely given the molecule volume, a crystal embedding model achieves R=0.76. In (c), and (d), as we increase the richness of the molecule representation, we see also the regression performance improves, as it is given access to the molecule shape and global orientation (principal vectors), and finally all the atom positions (autoencoder).
For the prediction of MACE MPA-0 lattice energies, we see the detailed information in the autoencoder embedding is much more important for accurate energy prediction. While the Buckingham potential is not very sensitive to the details of individual atom types and positions, the MACE MPA-0 potential, trained on density functional theory calculations, should be much more sensitive to the intricacies of particular atoms and environments. The principal vector embedding outperforms again the volume only or empty embeddings, but none give respectable regression results for the lattice energies. Only the autoencoder embedding with its detailed information on molecular position and pose, including the absolute spatial coordinates of all the atoms, is rich enough to achieve a reasonable accuracy.
In terms of training performance, we observed rapid saturation of the loss and minimal overfitting of the non-autoencoder embeddings due to the low resolution of any molecular features. The crystal embedding model trained on the autoencoder embedding, on the other hand, continues to improve with training time and increases in dataset size, and so we project its peak performance is even better than what we show in Figure 7. We are satisfied the given results make the desired point, and so we chose not to synthesize larger and larger datasets.
@@ 5 Discussion
We introduce and evaluate a new autoencoder architecture and reconstruction loss for the unsupervised learning of multi-type point clouds. This tool ‘solves’ the molecule representation problem in the limit of large data, in the sense that it can encode type, conformation, and pose information of a molecule in a provably complete embedding, via a sufficiently large and well trained autoencoder model. The practical limitations of our model are training time and dataset composition. To train Mo3ENet on a new class or distribution of molecules requires only sourcing or synthesizing a large and diverse pool of the relevant structures. The embedding quality on any unseen samples can be trivially evaluated by decoding the sample and computing the reconstruction loss.
Beyond high-fidelity reconstructions, we further proved the quality of the molecule embedding as an input for regressing the twenty QM9/QM9s molecule properties to high accuracy. We also developed a new type of crystal model, combining molecule embeddings with low-dimensional crystal parameter embeddings to predict properties of molecular crystal structures. We compared our autoencoder to simpler molecular embeddings and found the increased richness of its molecule representation, including atom types, molecular conformation and pose, was crucial for making accurate crystal energy predictions. Furthermore, the embedding is generally useful for downstream tasks where single molecules need to be represented in the context of larger material structures.
The autoencoder’s performance on molecule reconstruction, molecule property regression, and crystal modelling, combined with our qualitative embedding analyses, where we show visually that the model is clustering functional groups and structural motifs in its embedding space, indicate the model is not merely memorizing the input molecules, but generalizing across the space of molecular structures.
This work opens the door to several interesting applications. Such a pretrained molecule embedding will be valuable in future exploration of molecule clusters and materials. For example, in the context of molecular crystal structure prediction, this tool allows us to losslessly coarse-grain molecules to vector embeddings, while retaining the ability to rotate or invert the representations. This is potentially quite useful in the modelling of molecular clusters or crystals, as we have seen. This embedding is also potentially the linchpin in the inverse learning problem, i.e., the search, via a generative model, for low energy crystal parameters $\mathcal{C}$, for a given molecular conformer. Via training on larger and more diverse datasets, we can develop a universal/foundational representation model for high-fidelity encoding of molecule and pose information. Such a tool could be useful for molecule property prediction, similarity analysis/clustering, and conditioning for downstream learning tasks. The speed and generality of our approach also make it attractive as a new type of generative model. Finally, The freedom of the decoder to generate structures with arbitrary numbers, types, and positions of atoms, and the latent space’s amenability to conditioning, is tantalizing to explore further.